one school survey showed 3 out of 5 students own a pet. another survey showed that 6 out of 11 students own a pet. are these results proportional. explain your reasoning

Answer by fcabanski(1391)

A proportion means two ratios are equal. The ratio is student pet owners to students.


3 out of 5 is the ratio 3/5. 6 out of 11 is 6/11. Set them equal, then cross multiply. If the products are the same, then the ratios are equal and thus proportional.


3/5 = 6/11 --> 33 = 30? No. The ratios are not proportional.